Three times the sum of 8 and twice an unknown number is equal to 4 times the sum of 3 times the unknown number and 8. Find the number. I got 2.66, but it doesn't work. Also, what does "4 times the sum of 3 times the uknown number" mean mathmatically? : Three times the sum of 8 and twice an unknown number is equal to 4 times the sum of 3 times the unknown number and 8. Find the number. I got 2.66, but it doesn't work. Also, what does "4 times the sum of 3 times the uknown number" mean mathmatically?
Answer by jim_thompson5910(8291)

The statement "Three times the sum of 8 and twice an unknown number " translates to 3(8+2x) and the statement "4 times the sum of 3 times the unknown number and 8" translates to 4(3x+8). Altogether they form the equation 3(8+2x)=4(3x+8)


3(8+2x)=4(3x+8) Start with the given equation.


24+6x=12x+32 Distribute.


6x=12x+32-24 Subtract 24 from both sides.


6x-12x=32-24 Subtract 12x from both sides.


-6x=32-24 Combine like terms on the left side.


-6x=8 Combine like terms on the right side.


x=(8)/(-6) Divide both sides by -6 to isolate x.


x=-4/3 Reduce.


----------------------------------------------------------------------

Answer:

So the answer is x=-4/3


Which approximates to x=-1.333